In the aftermath of a hurricane, an entrepreneur took a one-month leave of absence (without pay) from her $5,000 per month job in order to operate a kiosk that sold fresh drinking water. During the moth she operated this venture, she paid the government $2,500 in kiosk rent and purchased water from a local wholesaler at a price of $1.34 per gallon. If consumers are willing to pay $2.25 to purchase each of gallon fresh drinking water, how many units will she have to sell in order to turn a profit? Provide your calculations and explain how this can be applied to other businesses. Why is this topic so important?
Answer #1

Implicit cost+ Explicit cost= $5000( loss of salary)+ $2500( kiosks rent)= $7500

Total variable cost= $1.34* Q

Cost function= C(Q)= 7500+1.34*Q

Total revenue= R(Q)= Price * Quantity= $2.25*Q

Profit= TR -TC= 2.25*Q-7500-1.34Q= $0.91Q-7500

Entrepreneur will only make profit when

0.91Q> 7500

Q> 8241.7

She have to 8241.7 atleast in order to trun a profit.

To apply this method to other businesses find the breakeven level of output, the output where total revenue is equal to total cost. So in order to earn positive profit, business has to sell more units than that. If that seems possible, the business couos be carried out else not.

This calculation is important because is allows entrepreneurs to make business decisions of whether to start or stay in a business or to shut down.